# Opening the Black Box of Local Projections

### v1

@@ 1 Introduction
Applicability. In the empirical analysis, we focus on OLS and Random Forest, the latter being a powerful, nonlinear, off-the-shelf machine learning algorithm. While not explicitly used here, the decomposition framework is fully compatible with any estimator that is, or can be rewritten as, a linear combination of outcomes. This includes kernel methods, boosting, and neural networks, drawing on tools developed in Goulet Coulombe et al., (2024). The same applies to two-stage least squares, frequently used alongside local projections. In that context, our tools could be applied separately to both the first and second stage regressions, with the second stage being conceptually closest to our empirical application. Finally, further thinking is required to extend the framework to impulse response estimators that are not linear aggregations of a single outcome variable—such as certain Bayesian methods or VARs, for which we sketch a possible route using Shapley values in the conclusion.
@@ 2 Historical Foundations of Local Projections
We briefly review LPs here and refer the reader to Jordà and Taylor, (2024) for a more detailed discussion. First, define $y_{t}$ as the outcome variable of interest at time $t$. The controls, denoted by $Z_{t}$, form a vector comprising an intercept, and exogenous or predetermined variables, including lagged values of both the outcome $y_{t}$ and the policy intervention $s_{t}$. Local projections provide a method to estimate the following population object. Formally, the impulse response is defined as
@@
where $\boldsymbol{X}\in\mathbb{R}^{T\times K}$ and $\boldsymbol{y}_{h}\in\mathbb{R}^{T}$ are matrices of training data. It is assumed that $s_{t}$ is placed in the second column of $\boldsymbol{X}=[\boldsymbol{1}\,\boldsymbol{s}\,\boldsymbol{Z}]$. By properties of projection and annihilator matrices, $\boldsymbol{1}^{\prime}\boldsymbol{w}=0$ because $\boldsymbol{X}$ includes an intercept. The expression of $\hat{\beta}_{h}$ as a weighted average of outcomes follows directly from the least squares formula and is noted, for instance, in Davidson and MacKinnon, (2004).111 Throughout the paper, all standard linear regression assumptions required for the consistency of point estimates are assumed to hold, as is customary in local projection applications. This includes coefficient stability, shock exogeneity, and inclusion of all relevant regressors. Importantly, these assumptions are not required for the decomposition itself to hold: the decomposition is purely mechanical and applies to any OLS estimate, regardless of its causal interpretation. What these assumptions ensure, rather, is that the object being decomposed is worth decomposing at all. Moreover, issues such as serial correlation and heteroskedasticity are typically addressed only at the inference stage, as they do not compromise consistency. Our decomposition of point estimates follows suit, reflecting how evidence is aggregated across the sample under OLS assumptions. Our key contribution is the introduction of $\boldsymbol{w}$ for interpretation purposes, along with its (eventual) alternative interpretation as proximity scores, which establishes a connection to machine learning methods that are linear in $\boldsymbol{y}_{h}$.
@@ 2.1 $w_{t}$ as a Standardized and Purified Shock Series
A numerically equivalent $\hat{\beta}_{h}$ can be obtained by running a regression where $y_{t+h}$ remains the target, and the sole regressor is $\tilde{s}_{t}$, the residual from regressing $s_{t}$ on all other variables in the original regression, $\boldsymbol{Z}_{t}$. Since this second regression is univariate, we have:
@@ 2.2 $w_{t}$ as a Proximity Weights Series
where $\tau$ is a hypothetical new observation, and $\boldsymbol{w}^{\mathcal{S}}=X_{\tau}^{\mathcal{S}}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}$, with $X_{\tau}^{\mathcal{S}}\equiv[\boldsymbol{1}\,\,\mathcal{S}\,\boldsymbol{Z}_{\tau}]$ for $\mathcal{S}\in\{0,\delta\}$, representing the two dosage scenarios. In Goulet Coulombe, (2025), OLS is reframed as a similarity-based estimator, where out-of-sample predictions are expressed as weighted linear combinations of the training values of the target variable, with weights interpreted as proximity scores. As shown in (9), local projection estimates can be computed—numerically equivalently—as the difference between two conditional forecasts at time $\tau$ for horizon $h$, given economic conditions $\boldsymbol{Z}_{\tau}$. [^3] In linear models, $\widehat{\text{IRF}}_{s\to y}(h,\delta)$ is invariant to the context $\boldsymbol{Z}_{\tau}$, rendering the IRF itself time-invariant. Thus, we can freely assign values to $\boldsymbol{Z}_{\tau}$ in linear models, as these contributions are canceled out in (9). However, retaining $\boldsymbol{Z}_{\tau}$ anyhow allows for an alternative interpretation of $\boldsymbol{w}$ as a measure of proximity, which becomes particularly meaningful when conditional expectations are nonlinear and the invariance to $\boldsymbol{Z}_{\tau}$ no longer holds. If the weight vectors $\boldsymbol{w}^{\delta}$ and $\boldsymbol{w}^{0}$ each carry a proximity interpretation, then so should their difference, $\boldsymbol{w}$.
@@ 2.2.1 A Proximity-Based Representation of $\hat{\beta}_{h}$: the Uncorrelated $X_{t}$ Case
$\hat{\beta}_{h}$ $=\frac{1}{T}\sum_{t=1}^{T}\left[\left\langle X_{\tau}^{\delta},X_{t}\right\rangle-\left\langle X_{\tau}^{0},X_{t}\right\rangle\right]y_{t+h}\,.$ (11)
Here, $w_{t}=\frac{\left\langle X_{\tau}^{\delta},X_{t}\right\rangle-\left\langle X_{\tau}^{0},X_{t}\right\rangle}{T}$, reflecting the proximity differential between conditions at time $t$ and the projected conditions for a dose of $\delta$ versus 0. In other words, $w_{t}$ increases if imposing $s_{\tau}=\delta$ on $X_{\tau}$ brings economic conditions at $\tau$ (represented as a point in $\mathbb{R}^{K}$) closer to those of $t$ than when setting $s_{\tau}=0$.
@@
$\hat{\beta}_{h}$ $=\frac{1}{T}\sum_{t=1}^{T}\left[\|{X}_{t}\|_{2}\left(\|X_{\tau}^{\delta}\|_{2}\cos(\theta^{\delta}_{\tau t})-\|X_{\tau}^{0}\|_{2}\cos(\theta^{0}_{\tau t})\right)\right]y_{t+h}.$ (12)
@@ 2.2.2 A Proximity-Based Representation of $\hat{\beta}_{h}$: the Correlated $X_{t}$ Case
$\hat{\beta}_{h}=\left[X_{\tau}^{\delta}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}-X_{\tau}^{0}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}\right]\boldsymbol{y}_{h},$
@@
$\hat{\beta}_{h}=\sum_{t=1}^{T}\left[\langle F_{\tau}^{\delta},F_{t}\rangle-\langle F_{\tau}^{0},F_{t}\rangle\right]y_{t+h},$
@@ 2.2.3 OLS as a Nearest-Interventions Estimator
$\hat{\beta}_{h}$ $=\frac{1}{T}\sum_{t=1}^{T}\left\langle X_{\tau}^{\delta}-X_{\tau}^{0},X_{t}-X_{t}^{0}\right\rangle y_{t+h},$ (13)
@@
$\hat{\beta}_{h}$ $=\sum_{t=1}^{T}\left\langle F_{\tau}^{\delta}-F_{\tau}^{0},F_{t}-\boldsymbol{0}\right\rangle y_{t+h}.$ (14)
@@
$\hat{\beta}_{h}$ $=\sum_{t=1}^{T}\left\langle\mathcal{I}_{\tau},\mathcal{I}_{t}\right\rangle y_{t+h}=\sum_{t=1}^{T}\cos(\theta^{\mathcal{I}}_{\tau t})\|\mathcal{I}_{\tau}\|_{2}\|\mathcal{I}_{t}\|_{2}\,y_{t+h}.$ (15)
Here, $w_{t}$ corresponds to the inner product proximity scores between $\mathcal{I}_{\tau}$ and $\mathcal{I}_{t}$, which are not scale-invariant and embed information about both vector alignment and magnitudes. The second representation, based on cosine similarity, explicitly separates these components: $w_{t}$ is expressed as the product of the scale-invariant alignment in $\mathbb{R}^{K}$ (cosine similarity) and the magnitudes (measured by the $l_{2}$-norm) of the two interventions being compared. Intuitively, a large $\|\mathcal{I}_{\tau}\|_{2}$ (e.g., due to a high $\delta$) scales $\hat{\beta}_{h}$ linearly upward. A large $\|\mathcal{I}_{t}\|_{2}$ indicates that observation $t$ is particularly informative and carries leverage, being located farther from the origin $\boldsymbol{0}$ in the orthonormal feature space. In Section 3, we report the two components separately ($\cos(\theta^{\mathcal{I}}_{\tau t})$ and $|\mathcal{I}_{t}|_{2}$) to disentangle the respective roles of alignment and scale in the baseline monetary policy results.
@@ 2.3.2 Retrieving Weights in Random Forest
for intervention scenario $\mathcal{S}$, context indexed by $\tau$, and the horizon $h$. In words, to generate $\boldsymbol{w}_{\tau h}^{\mathcal{S}}$ in the RF case, one can follow these steps: determine which leaf observation $\tau$ falls into for a given tree (based on its $X_{\tau}^{\mathcal{S}}$), identify the corresponding in-sample observations for the leaf and their weights (calculated as 1/leaf $\nicefrac{{1}}{{\text{leaf size}}}$), assign these weights to the relevant in-sample observations ($w_{b\tau th}$), and then aggregate these “votes” across all trees in the ensemble.
@@ 3.1 Monetary Policy
The main specification includes inflation, unemployment, industrial production, federal funds rate, S&P500 stock market index, and the respective shock series. The latter is obtained from i) R&R and ii) a Cholesky VAR estimated with the listed variables (excluding the shock) in the same order. Data is taken from FRED-MD (McCracken and Ng,, 2016), with all variables including 12 lags. We report cumulative IRFs throughout, including 84% confidence bands. To account for serial correlation in the error term, we apply the Newey-West correction to the standard errors (Newey and West,, 1987).444 If a less passive approach to heteroskedasticity—such as stochastic volatility or generalized least squares (GLS)—were used, the contributions and weights would adjust accordingly. This raises an intriguing possibility: although feasible GLS is often set aside in favor of OLS with corrected inference, visualizing the implied data weighting may help users better understand and accept differences in point estimates.
@@
Nonlinear Enlightenment. We complete the analysis with our nonlinear IRFs for R&R shocks (see Figure 4), addressing two questions: First, is there any evidence of effects from tightening shocks in an R&R shock-based local projection? Second, does the early hump observed with R&R shocks persist in a nonlinear model where shock size and sign matter? The answer to the first question is negative—IRFs from R&R tightening shocks are effectively null. R&R acknowledge in their original paper that much of the exogenous variation in interest rates is attributable to political influence—and that the pressure is almost always unidirectional. Therefore, it is natural to ask whether there is any quantitative evidence from R&R we should use to forecast the effects of future monetary tightening decisions.
@@ 3.2 Government Spending
In Figure 20 (Appendix), we see that both shocks yield relatively similar estimates for $h=5$, with $\hat{\beta}_{5}$ heavily influenced by the Korean War, accounting for more than half of the reported value in both cases. However, results diverge at $h=9$ and $h=12$, where BZP produces more moderate estimates compared to $h=5$, while the IRF from Ramey, (2011) remains elevated and even increases further. This discrepancy arises partly from how the years immediately preceding the Korean War contribute to later horizons. For Ramey, (2011), these years contribute in the same direction as earlier horizons, whereas for BZP, they push the IRF toward zero. There is some indication that these estimates might remain viable even without their most influential historical contribution, as the curves steadily climb in the expected direction from the 1950s onward. Additionally, concentration measures suggest that the distribution of weights as well as concentrations are more dispersed (with a WC($\hat{\beta}_{h}$) of 40% and CC($\hat{\beta}_{h}$) ranging from 0.52 and 0.58 across different horizons).
@@ 3.3 Global Temperature Shocks
Results are presented in Figures 7 and 8, which display cumulative contributions and the corresponding weights alongside the values of the response variable, respectively. By examining the evidence curves derived from the key impulse response function for world GDP, we observe the clear influence of two significant climatological events on the estimated effects. These events are the 1964 cooling shock, the late-1990s El Niño, and the subsequent La Niña event. Despite these observations, the overall concentration levels are more moderate than in previous applications, with WC$(\hat{\beta}_{h})=0.25$ and CC($\hat{\beta}_{h}$) ranging from 0.27 ($h=3$) to 0.33 ($h=10$). Therefore, some more investigation is needed.

### v2

@@ 1 Introduction
Applicability. In the empirical analysis, we focus on OLS and Random Forest, the latter being a powerful, nonlinear, off-the-shelf machine learning algorithm. While not explicitly used here, the decomposition framework is fully compatible with any estimator that is, or can be rewritten as, a linear combination of outcomes. This includes kernel methods, boosting, and neural networks, drawing on tools developed in Goulet Coulombe et al., (2024). The same applies to two-stage least squares, frequently used alongside local projections. In that context, our tools could be applied separately to both the first and second stage regressions, with the second stage being conceptually closest to our empirical application. The tool could also be applied in a panel data setting to decompose how both time periods and individual units contribute to the estimated effects. Finally, further thinking is required to extend the framework to impulse response estimators that are not linear aggregations of a single outcome variable—such as certain Bayesian methods or VARs, for which we sketch a possible route using Shapley values in the conclusion.
@@ 2 Historical Foundations of Local Projections
We briefly review LPs here and refer the reader to Jordà and Taylor, (2024) for a more detailed discussion. First, define $y_{t}$ as the outcome variable of interest at time $t$. The controls, denoted by $Z_{t}$, form a vector comprising exogenous or predetermined variables, including lagged values of both the outcome $y_{t}$ and the policy intervention $s_{t}$. Local projections provide a method to estimate the following population object. Formally, the impulse response is defined as
@@
[…]
@@ 2.1 $w_{t}$ as a Standardized and Purified Shock Series
A numerically equivalent $\hat{\beta}_{h}$ can be obtained by running a regression where $y_{t+h}$ remains the target, and the sole regressor is $\tilde{s}_{t}$, the residual from regressing $s_{t}$ on all other variables in the original regression, ${Z}_{t}$. Since this second regression is univariate, we have:
@@ 2.2 $w_{t}$ as a Proximity Weights Series
where $\tau$ is a hypothetical new observation, and $\boldsymbol{w}^{\mathcal{S}}=X_{\tau}^{\mathcal{S}}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}$, with $X_{\tau}^{\mathcal{S}}\equiv[{1}\,\,\mathcal{S}\,{Z}_{\tau}]\in\mathbb{R}^{K}$ for $\mathcal{S}\in\{0,\delta\}$, representing the two dosage scenarios. In Goulet Coulombe, (2025), OLS is reframed as a similarity-based estimator, where out-of-sample predictions are expressed as weighted linear combinations of the training values of the target variable, with weights interpreted as proximity scores. As shown in (9), local projection estimates can be computed—numerically equivalently—as the difference between two conditional forecasts at time $\tau$ for horizon $h$, given economic conditions ${Z}_{\tau}$. [^4] In linear models, $\widehat{\text{IRF}}_{s\to y}(h,\delta)$ is invariant to the context ${Z}_{\tau}$, rendering the IRF itself time-invariant. Thus, we can freely assign values to ${Z}_{\tau}$ in linear models, as these contributions are canceled out in (9). However, retaining ${Z}_{\tau}$ anyhow allows for an alternative interpretation of $\boldsymbol{w}$ as a measure of proximity, which becomes particularly meaningful when conditional expectations are nonlinear and the invariance to ${Z}_{\tau}$ no longer holds. If the weight vectors $\boldsymbol{w}^{\delta}$ and $\boldsymbol{w}^{0}$ each carry a proximity interpretation, then so should their difference, $\boldsymbol{w}$.
@@ 2.2.1 A Proximity-Based Representation of $\hat{\beta}_{h}$: the Uncorrelated $X_{t}$ Case
$\hat{\beta}_{h}$ $=\frac{1}{\delta T}\sum_{t=1}^{T}\left[\left\langle X_{\tau}^{\delta},X_{t}\right\rangle-\left\langle X_{\tau}^{0},X_{t}\right\rangle\right]y_{t+h}\,.$ (11)
Here, $w_{t}=\frac{\left\langle X_{\tau}^{\delta},X_{t}\right\rangle-\left\langle X_{\tau}^{0},X_{t}\right\rangle}{\delta T}$, reflecting the proximity differential between conditions at time $t$ and the projected conditions for a dose of $\delta$ versus 0. In other words, $w_{t}$ increases if imposing $s_{\tau}=\delta$ on $X_{\tau}^{\mathcal{S}}$ brings economic conditions at $\tau$ (represented as a point in $\mathbb{R}^{K}$) closer to those of $t$ than when setting $s_{\tau}=0$. We reintroduce the more general dosage parameter $\delta$ in place of 1. Since the proximity interpretation extends to machine learning methods, where the magnitude and sign of $\delta$ affect impulse response estimates, retaining $\delta$—though redundant in linear models—is preferable.
@@
$\hat{\beta}_{h}$ $=\frac{1}{\delta T}\sum_{t=1}^{T}\left[\|{X}_{t}\|_{2}\left(\|X_{\tau}^{\delta}\|_{2}\cos(\theta^{\delta}_{\tau t})-\|X_{\tau}^{0}\|_{2}\cos(\theta^{0}_{\tau t})\right)\right]y_{t+h}.$ (12)
@@ 2.2.2 A Proximity-Based Representation of $\hat{\beta}_{h}$: the Correlated $X_{t}$ Case
$\hat{\beta}_{h}=\frac{1}{\delta}\left[X_{\tau}^{\delta}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}-X_{\tau}^{0}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}\right]\boldsymbol{y}_{h},$
@@
$\hat{\beta}_{h}=\frac{1}{\delta}\sum_{t=1}^{T}\left[\langle F_{\tau}^{\delta},F_{t}\rangle-\langle F_{\tau}^{0},F_{t}\rangle\right]y_{t+h},$
@@ 2.2.3 OLS as a Nearest-Interventions Estimator
$\hat{\beta}_{h}$ $=\frac{1}{\delta T}\sum_{t=1}^{T}\left\langle X_{\tau}^{\delta}-X_{\tau}^{0},X_{t}-X_{t}^{0}\right\rangle y_{t+h},$ (13)
@@
$\hat{\beta}_{h}$ $=\frac{1}{\delta}\sum_{t=1}^{T}\left\langle F_{\tau}^{\delta}-F_{\tau}^{0},F_{t}-\boldsymbol{0}\right\rangle y_{t+h}.$ (14)
@@
$\hat{\beta}_{h}$ $=\frac{1}{\delta}\sum_{t=1}^{T}\left\langle\mathcal{I}_{\tau},\mathcal{I}_{t}\right\rangle y_{t+h}=\frac{\|\mathcal{I}_{\tau}\|_{2}}{\delta}\sum_{t=1}^{T}\cos(\theta^{\mathcal{I}}_{\tau t})\|\mathcal{I}_{t}\|_{2}\,y_{t+h}.$ (15)
Here, $w_{t}$ corresponds to the inner product proximity scores between $\mathcal{I}_{\tau}$ and $\mathcal{I}_{t}$, which are not scale-invariant and embed information about both vector alignment and magnitudes. The second representation, based on cosine similarity, explicitly separates these components: $w_{t}$ is expressed as the product of the scale-invariant alignment in $\mathbb{R}^{K}$ (cosine similarity) and the magnitudes (measured by the $l_{2}$-norm) of the two interventions being compared. Intuitively, a large $\|\mathcal{I}_{\tau}\|_{2}$ (e.g., due to a high $\delta$) scales $\hat{\beta}_{h}$ linearly upward. A large $\|\mathcal{I}_{t}\|_{2}$ indicates that observation $t$ is particularly informative and carries leverage, being located farther from the origin $\boldsymbol{0}$ in the orthonormal feature space.555 In the uncorrelated case, $\|\mathcal{I}_{\tau}\|_{2}=|\delta|$, so the expression simplifies to $\hat{\beta}_{h}=\text{sign}(\delta)\times\sum_{t=1}^{T}\cos(\theta^{\mathcal{I}}_{\tau t})\|\mathcal{I}_{t}\|_{2}\,y_{t+h}$. In Section 3, we report the two components separately ($\cos(\theta^{\mathcal{I}}_{\tau t})$ and $|\mathcal{I}_{t}|_{2}$) to disentangle the respective roles of alignment and scale in the baseline monetary policy results.
@@ 2.3.2 Retrieving Weights in Random Forest
for intervention scenario $\mathcal{S}$, context indexed by $\tau$, and the horizon $h$. In words, to generate $\boldsymbol{w}_{\tau h}^{\mathcal{S}}$ in the RF case, one can follow these steps: determine which leaf observation $\tau$ falls into for a given tree (based on its $X_{\tau}^{\mathcal{S}}$), identify the corresponding in-sample observations for the leaf and their weights (calculated as $\nicefrac{{1}}{{\text{leaf size}}}$), assign these weights to the relevant in-sample observations ($w_{b\tau th}$), and then aggregate these “votes” across all trees in the ensemble.
@@ 3.1 Monetary Policy
The main specification includes inflation, unemployment, industrial production, federal funds rate, S&P500 stock market index, and the respective shock series. The latter is obtained from i) R&R and ii) a Cholesky VAR estimated with the listed variables (excluding the shock) in the same order. Data is taken from FRED-MD (McCracken and Ng,, 2016), with all variables including 12 lags. We report cumulative IRFs throughout, including 84% confidence bands. To account for serial correlation in the error term, we apply the Newey-West correction to the standard errors (Newey and West,, 1987).666 If a less passive approach to heteroskedasticity—such as stochastic volatility or generalized least squares (GLS)—were used, the contributions and weights would adjust accordingly. This raises an intriguing possibility: although feasible GLS is often set aside in favor of OLS with corrected inference, visualizing the implied data weighting may help users better understand and accept differences in point estimates.
@@
Nonlinear Enlightenment. We complete the analysis with our nonlinear IRFs for R&R shocks (see Figure 4). The nonlinear model is a Random Forest, following the methodology outlined in Section 2.3, with implementation details provided in Appendix A.2.
We address two questions: First, is there any evidence of effects from tightening shocks in an R&R shock-based local projection? Second, does the early hump observed with R&R shocks persist in a nonlinear model where shock size and sign matter? The answer to the first question is negative—IRFs from R&R tightening shocks are effectively null. R&R acknowledge in their original paper that much of the exogenous variation in interest rates is attributable to political influence—and that the pressure is almost always unidirectional. Therefore, it is natural to ask whether there is any quantitative evidence from R&R we should use to forecast the effects of future monetary tightening decisions.
@@ 3.2 Government Spending
[…]
@@ 3.3 Global Temperature Shocks
[…]